A bag contains three white balls numbered 2, 4,6, and flour black balls numbered 1,3,5 and 7, two balls are drawn without replacement.
what is the probability the first ball is white given that the second ball is black?

Answer by greenestamps(13198)



First ball drawn: There are 7 balls in all; 3 are white. P(white on first draw) = 3/7.

Second ball drawn: Now there are 6 balls in all; 4 are black. P(black on second draw) = 4/6 = 2/3.

P(white on first AND black on second) = P(white on first) * P(black on second) = (3/7)*(2/3) = 2/7.